Jane leaves her house and drives in a straight line at a speed of 30 mph. Her speed on the return trip is 40 mph. The trip back home takes 10 minutes less than the outbound trip. What was Jane's total travel time in hours? Assume both trips are in a straight line. Give your final answer as a simplified fraction.

Answer by richwmiller(17219)

r*t=d
30*t=d,
40*(t-1/6)=d
30*t=d,
40*(t-1/6)=30*t
40*t-1/6*40=30*t
10*t=1/6*40
divide by 10
t=4/6=2/3
BTW d=20